OpenStudy (anonymous):

How do you write 24x+6y=-60 in slope intercept

OpenStudy (amistre64):

"solve" for y

OpenStudy (anonymous):

The slope intercept form is y =mx+c \[24*x+6*y=-60\] It can be written as \[6y = -24x - 60\] or y = -4x -10